OpenStudy (anonymous):

What is the equation of the line passing through (2, –3) and parallel to the line y = –6x – 1 in standard form?

OpenStudy (zzr0ck3r):

are you just pasting every homeowrk problem here? I dont think this will help. This is pretty much like the last 4

OpenStudy (zzr0ck3r):

what is standard form?

OpenStudy (zzr0ck3r):

y = mx+b and you know m so use the point to find b

OpenStudy (zzr0ck3r):

@kbreezy102 ?

OpenStudy (anonymous):

4000

OpenStudy (zzr0ck3r):

if y = -3, x = 2, and m = -6 what is b?

OpenStudy (anonymous):

Parallel means that the line will have the same slope as the reference equation. Standard form is, by definition, Ax + By = C. The original equation is in the form y = mx + b, so we convert it. y = 6x-1 y - 6x = -1 Now, the new line passes through (2, -3). We can write this in point-slope form to make it easier and convert to standard form: y1 - y0 = m(x1- x0) y +3 = 6(x-2) (because subtracting from a negative is addition) Now convert to standard form: y+3=6x - 12 y - 6x + 3 = -12 y - 6x = -15

OpenStudy (zzr0ck3r):

I feel we are wasting our time

OpenStudy (anonymous):

Lol maybe ...